OpenStudy (anonymous):

(2x+3x-6) + (x-18+6)=

OpenStudy (anonymous):

what are the rules? to solving this problem?

OpenStudy (anonymous):

Pretend the parentheses arent there, and combine like terms. Do you know how to do that?

OpenStudy (anonymous):

find the common factor?

OpenStudy (anonymous):

i mean the lowest common factor

OpenStudy (anonymous):

No, combine like terms

OpenStudy (anonymous):

2x+x=3x?

OpenStudy (anonymous):

2x+3x=5x-6

OpenStudy (anonymous):

Combining like terms: Combine (add or subtract) everything that has the same base/ is the same. for an example, if there is 6x+2, you cannot combine that because 2 does not have an x term with it. However, you can do 3x+7x (=10x) because they both have a variable attached (x). You can also do a simple 5+1(=6) because they both have the same base

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

So based on that, how can you simplify the first set of parentheses?

OpenStudy (anonymous):

5x-6

OpenStudy (anonymous):

Right! Now the second one?

OpenStudy (anonymous):

the second is x-18x+6

OpenStudy (anonymous):

Oh so there is an x with the 18?

OpenStudy (anonymous):

17x

OpenStudy (anonymous):

Not really.. whats 1-18?

OpenStudy (anonymous):

17x+6

OpenStudy (anonymous):

-17x+6

OpenStudy (anonymous):

Now thats better.. 5x-6-17x+6=0 Now combine that

OpenStudy (anonymous):

-12x=0

OpenStudy (anonymous):

or -12x/-12 =0/-12

OpenStudy (anonymous):

Right, so x=0

OpenStudy (anonymous):

yay!!!! thank you

OpenStudy (anonymous):

No problem!